in the x-y plane, the equation y=1/2x-3 is a line that is perpendicualr to the graph of the equation y=mx+(1/3). what is the value of m?
Found 2 solutions by Alan3354, Maths68:
Answer by Alan3354(69443)

m = the negative inverse of 1/2
m = -2

Answer by Maths68(1474)

y=1/2x-3 ...........(1)
y=mx+(1/3)...........(2)
Compare (1) with the Equation of the straight line standard form
y=1/2x-3 ...........(1)
y=mx+b
m=(1/2), b=-3
When two lines are perpendicular the multiplication of the their slopes result in (-1)
So the slope of the other line will be = m = -2